# DOCTOR: Dynamic On-Chip Remediation Against Temporally-Drifting Thermal Variations Toward Self-Corrected Photonic Tensor Accelerators

### v1

@@ Abstract
Photonic computing has emerged as a promising solution for accelerating computation-intensive artificial intelligence (AI) workloads, offering unparalleled speed and energy efficiency, especially in resource-limited, latency-sensitive edge computing environments. However, the deployment of analog photonic tensor accelerators encounters reliability challenges due to hardware noises and environmental variations. While off-chip noise-aware training and on-chip training have been proposed to enhance the variation tolerance of optical neural accelerators with moderate, static noises, we observe a notable performance degradation over time due to temporally drifting variations, which requires a real-time, in-situ calibration mechanism. To tackle this challenging reliability issues, for the first time, we propose a lightweight dynamic on-chip remediation framework, dubbed DOCTOR, providing adaptive, in-situ accuracy recovery against temporally drifting noises. The DOCTOR framework intelligently monitors the chip status using adaptive probing and performs fast in-situ training-free calibration to restore accuracy when necessary. Recognizing nonuniform spatial variation distributions across devices and tensor cores, we also propose a variation-aware architectural remapping strategy to avoid executing critical tasks on noisy devices. Extensive experiments show that our proposed framework can guarantee sustained performance under drifting variations with 34% higher accuracy and 2-3 orders-of-magnitude lower overhead compared to state-of-the-art on-chip training methods.
@@ I Introduction
However, the deployment of such accelerators encounters robustness challenges that impede their practical application [10, 11, 12, 13]. We consider one of the most sensitive accelerators based on micro-ring resonators (MRRs) as a case study [14, 15, 9]. Due to the intrinsic temperature sensitivity of the MRR device, shown in Fig. 1(a), a subtle drift in the temperature will lead to a slight change of the round-trip phase shift but a large deviation on the represented weight. Such a high sensitivity makes pre-deployment optimization ineffective and thus necessitates a real-time calibration mechanism on chip. Besides temperature drift, various dynamic random noises and crosstalk cast even more shadows on the reliability of photonic computing systems. Figure 1(b) shows the significant impacts of variations on accuracy, sometimes leading to malfunction over time when the noise intensities gradually increase.
While previous off-chip noise-aware model training [10, 16] have shown efficacy in enhancing the variation tolerance of optical accelerators by injecting noises during training and thus encourage a smoother solution space, they rely on accurate noise modeling and thus typically show unsatisfying robustness improvement with unknown physical variations and can only handle small and static noises. The performance drop remains unresolved when there exist temporally drifting variations. Recently, there has been a trend to resort to on-chip learning or physical training methods to directly train the optical neural network (ONN) models in situ that can naturally incorporate real physical noises into the weight training process [17, 18, 19, 20, 21]. However, they require repeatedly performing forward and backward propagation of the entire network on a labeled training dataset to calculate the task-specific gradients for weight fine-tuning, which induces nontrivial training costs that can severely harm the system throughput and efficiency. Moreover, prior methods fail to leverage the nonuniform spatial noise distribution and weight sensitivity, shown in Fig. 1(c), to balance accuracy and efficiency. Hence, it necessitates a real-time, low-cost, in-situ calibration mechanism without running backpropagation on any labeled training set to quickly recover the accuracy and ensure continued reliability in practical deployment.
@@
Figure 1: (a) MRR-based photonic accelerator is sensitive to temperature drift. (b) Drifting noises cause severe accuracy drop over time, including phase variation (PV), temperature drift (TD), and thermal crosstalk (CT). (c) Noise distributions across devices are nonuniform.
@@
    Salience-Aware Sparse Calibration: We propose a training-free, data-free in-situ calibration mechanism to selectively cancel out the device thermal variations and efficiently resume accuracy with negligible overhead.
@@
    Variation-Aware Tile Remapping: We leverage the nonuniform spatial noise distributions to boost the reliability by optimally remapping workloads onto tensor cores, aware of the weight importance and device noise levels.
@@
    We evaluate that our DOCTOR framework guarantees sustained deployment performance with 1%-2.5% accuracy drop at negligible runtime overhead (0.1%-5% cycle overhead), outperforming state-of-the-art on-chip training methods by an average of 34% higher accuracy and 2-3 orders-of-magnitude less overhead. Our work makes significant strides toward the real-world deployment of photonic accelerators in dynamic environments.
@@ II-A Photonic Tensor Accelerators
Various photonic neural network designs have been proposed and demonstrated to encode inputs and weights to the light magnitude/phase and circuit transmission, respectively, and perform ultra-fast matrix multiplication [1, 2, 3, 4, 5]. Typically, the photonic circuits are sensitive to thermal variation as temperature impacts the refractive index of the optical component. Especially for compact microring resonator (MRR)-based photonic accelerators, the weight $w$ is encoded by the transmission $a$ of the add-drop MRR as $w=g(2a-1);~{}a=\frac{\alpha^{2}-2r\alpha\cos{\phi}+r^{2}}{1-2r\alpha\cos{\phi}+r^{2}\alpha^{2}}\in(0,1)$, where $\alpha$ is the attenuation factor, $r$ is the coupling factor, $\phi$ is the round-trip phase shift, and $g$ is the scaling factor. To accommodate multiple wavelengths in a single free-spectral range (FSR), the transmission of the MRR is designed with a sharp curve. This design choice makes the MRR highly sensitive to small disturbances, highlighting the urgency in addressing challenges related to thermal variation robustness.
@@ II-B Noise-Aware ONN Optimization
Noise-aware optimization for photonic accelerators includes two categories, i.e., offline optimization and on-chip optimization. Prior offline methods inject noises into ONN model training to obtain smooth solution space for better noise tolerance [10]. However, such a method requires accurate noise modeling and can only handle static, known variations, which cannot be adapted to dynamically drifting, unknown variations. On-chip training, as a trend, has been demonstrated for online adaptability and efficiency in in-situ accuracy recovery. The pretrained optical NNs are fine-tuned on a target dataset with on-chip gradient calculation [17, 18, 19]. However, they rely on accurate gradient calculation and require costly forward and backward propagation on a given training set, which can significantly degrade the edge inference throughput and efficiency and bear data privacy issues.
@@ III-B1 Device Phase Variation
Beside nonuniform noise distribution, the noise profiles are dynamically drifting along time, as shown in Fig. 5. We simulate such dynamics with two-level sampling. For each PTC, we use the std ($\sigma_{ij}$) of the phase noise distribution to represent the noise intensity for the MRR at coordinate $ij$, while the intensity itself is time varying. At time step $t$, we sample the step-$t$ noise intensity from a distribution, i.e., $\sigma_{ij}^{t+1^{\prime}}\sim\mathcal{N}(\mu_{s}(t),\sigma_{s}^{2}(t))$. We apply exponential moving average $\sigma_{ij}^{t+1}\leftarrow\beta_{\sigma}\sigma_{ij}^{t}+(1-\beta_{\sigma})\sigma_{ij}^{t+1^{\prime}}$ with a dampling factor $\beta_{\sigma}$ to smooth out the intensity drifting. $\mu_{s}(t)$ and $\sigma_{s}(t)$ are temporally evolving with a certain scheduling function.
@@ III-B2 Environmental Temperature Drift
The environmental temperature fluctuation can impact the reliability of the chip, especially for thermal-sensitive microring resonator-based PTCs [26, 22, 23]. As shown in Fig. 1(b), for MRRs, even 0.05K temperature drift can lead to a relatively large resonant wavelength shift, leading to large errors on the represented weight value. We assume a linear dependence of on-resonance wavelength $\lambda_{c}$ on temperature $T$, thus, we assume a constant $\partial\lambda/\partial T$. The drift on the round-trip phase shift $\delta\phi$ due to temperature change is derived as,
@@ III-C Salience-Aware Sparse Calibration
(2) Accurate gradients – Unlike on-chip training, where gradient errors will exponentially accumulate through layers, our method only calculates gradients for each block without backpropagation. Hence, our method avoids slow convergence or divergence issues even under large noises.
@@ IV-A Simulation Setup
Dataset and Models.  We evaluate our method on a three-layer CNN (C64K3-C64K3-C64K3-Pool5-FC10) on Fashion-MNIST, VGG-8 on CIFAR-10, and ResNet-18 CIFAR-100 for image classification.
@@
Architecture Settings.  As a challenging case study, the hardware platform in this work is assumed to be a multi-tile, multi-core photonic tensor accelerator based on thermally sensitive MRR weight bank shown in Fig. 3. Note that our method is not specific to MRR weight banks but can generalize to all universal optical matrix-vector multiplication units. We assume that the photonic accelerator has 4 tiles, and each tile has 4 cores. Each core is an 8$\times$8 add-drop MRR weight bank that can perform 8$\times$8 matrix-vector multiplication per core per cycle. Detailed architecture description is in Section III-A. The MRR device modeling is based on Section II-A. The device/circuit variation modeling is based on Section III-B.
@@
Phase Variation (PV): To simulate real-world chip noise scenarios, two cases are created: "Low Noise" and "High Noise," corresponding to low and high standard deviation (std.) in phase shift, modeled as $\Delta\phi\in\mathcal{N}(0,\sigma^{2})$. As explained in Section III-B, the noise std. $\sigma_{ij}^{t+1^{\prime}}$ for each MRR device is time-variant and sampled at each time step from a noise level map $\mathcal{N}\in(\mu_{s}(t),\sigma_{s}^{2}(t))$. The noise level distribution gradually changes its high-noise region from the chip’s left edge to the top-left corner ("Edge-to-Corner") to capture dynamic noise profile drifting. Specific noise level functions are chosen for the Low and High Noise cases:
@@
Temperature Drift (TD): Two cases are designed to represent different types of temperature change with $t_{max}=20,000$:
@@
Thermal Crosstalk (CT):, we use a default MRR spacing $l_{v}=200\mu m$, $l_{h}=60\mu m$, crosstalk coefficient $k_{1}=0.1$, and model crosstalk only among MRRs within the same PTC [27, 24, 29].
@@ IV-C Main Results: Compare with Prior Work
Across all benchmarks, BS1 suffers from severe accuracy degradation (30%-60%). Noise-aware training [10], though it enhances the smoothness of the solution space of the model, shows limited effect on accuracy improvement, as the noise distribution used in pre-training is significantly different from the physical variations. Note that our DOCTOR is orthogonal to noise-aware training, allowing simultaneous application for a solution space that is both locally smooth and adaptive to the drifting noise distribution. On-chip training [19, 21] can boost the accuracy on small networks, e.g., 8% on CNN-FMNIST and VGG8-CIFAR10, however, it performs poorly on deep ResNet18 and not stable on certain benchmarks, e.g., CT+PV.2+TD.1. The fundamental reason is that the gradient estimation error exponentially accumulates with backpropagation, which can lead to poor training performance and even divergence. Also, though it is only trained for 1 epoch on a 10% training set, it consumes 5-6$\times$ more runtime (cycles) in training, which is not practical to be deployed on throughput-restricted edge platforms. In contrast, our proposed DOCTOR method can stably maintain high accuracy even with rapidly drifting noise distributions and temperature with less than 1-2.5% accuracy drop at the cost of merely 0.1%-5.1% cycle overhead. On average, DOCTOR is +34% more accurate and 2-3 orders-of-magnitude more efficient than on-chip training. We visualize our DOCTOR flow in Fig. 13 to show the temperature drift is detected and the accuracy is rapidly resumed with only 3.6% cycle overhead.

### v2

@@ Abstract
Photonic computing has emerged as a promising solution for accelerating computation-intensive artificial intelligence (AI) workloads, offering unparalleled speed and energy efficiency, especially in resource-limited, latency-sensitive edge computing environments. However, the deployment of analog photonic tensor accelerators encounters reliability challenges due to hardware noise and environmental variations. While off-chip noise-aware training and on-chip training have been proposed to enhance the variation tolerance of optical neural accelerators with moderate, static noise, we observe a notable performance degradation over time due to temporally drifting variations, which requires a real-time, in-situ calibration mechanism. To tackle this challenging reliability issues, for the first time, we propose a lightweight dynamic on-chip remediation framework, dubbed DOCTOR, providing adaptive, in-situ accuracy recovery against temporally drifting noise. The DOCTOR framework intelligently monitors the chip status using adaptive probing and performs fast in-situ training-free calibration to restore accuracy when necessary. Recognizing nonuniform spatial variation distributions across devices and tensor cores, we also propose a variation-aware architectural remapping strategy to avoid executing critical tasks on noisy devices. Extensive experiments show that our proposed framework can guarantee sustained performance under drifting variations with 34% higher accuracy and 2-3 orders-of-magnitude lower overhead compared to state-of-the-art on-chip training methods. Our code is open-sourced at link.
@@ I Introduction
However, the deployment of such accelerators encounters robustness challenges that impede their practical application [10, 11, 12, 13]. We consider one of the most sensitive accelerators based on micro-ring resonators (MRRs) as a case study [14, 15, 9]. Due to the intrinsic temperature sensitivity of the MRR device, shown in Fig. 1(a), a subtle drift in the temperature will lead to a slight change of the round-trip phase shift but a large deviation on the represented weight. Such a high sensitivity makes pre-deployment optimization ineffective and thus necessitates a real-time calibration mechanism on chip. Besides temperature drift, various dynamic random noise and crosstalk cast even more shadows on the reliability of photonic computing systems. Figure 1(b) shows the significant impacts of variations on accuracy, sometimes leading to malfunction over time when the noise intensities gradually increase.
While previous off-chip noise-aware model training [10, 16] have shown efficacy in enhancing the variation tolerance of optical accelerators by injecting noise during training and thus encourage a smoother solution space, they rely on accurate noise modeling and thus typically show unsatisfying robustness improvement with unknown physical variations and can only handle small and static noise [10, 11, 17, 18]. The performance drop remains unresolved when there exist temporally drifting variations. Recently, there has been a trend to resort to on-chip learning or physical training methods to directly train the optical neural network (ONN) models in situ that can naturally incorporate real physical noise into the weight training process [19, 20, 21, 18, 22]. However, they require repeatedly performing forward and backward propagation of the entire network on a labeled training dataset to calculate the task-specific gradients for weight fine-tuning, which induces nontrivial training costs that can severely harm the system throughput and efficiency. Moreover, prior methods fail to leverage the nonuniform spatial noise distribution and weight sensitivity, shown in Fig. 1(c), to balance accuracy and efficiency. Hence, it necessitates a real-time, low-cost, in-situ calibration mechanism without running backpropagation on any labeled training set to quickly recover the accuracy and ensure continued reliability in practical deployment.
@@
Figure 1: (a) MRR-based photonic accelerator is sensitive to temperature drift. (b) Drifting noises cause severe accuracy drop over time, including phase variation (PV), temperature drift (TD), and thermal crosstalk (CT). "Acc" represents the accuracy. (c) Noise distributions across devices are nonuniform.
@@
    Salience-Aware Sparse Calibration: We propose a training-free, data-free in-situ calibration mechanism to selectively mitigate thermal variations and effectively resume computing accuracy at negligible runtime overhead.
@@
    Variation-Aware Tile Remapping: We leverage the spatial nonuniformity in noise distributions to boost the reliability by optimally remapping workloads onto tensor cores, aware of the weight importance and device noise levels.
@@
    We evaluate that our DOCTOR framework guarantees sustained deployment performance with 1%-2.5% accuracy drop at negligible runtime overhead (0.1%-5% cycle overhead), outperforming state-of-the-art on-chip training methods [21, 22] by an average of 34% higher accuracy and 2-3 orders-of-magnitude less overhead. Our work makes significant strides toward the real-world deployment of photonic accelerators in dynamic environments.
@@ II-A Photonic Tensor Accelerators
Various photonic neural network designs have been proposed and demonstrated to encode inputs and weights to the light magnitude/phase and circuit transmission, respectively, and perform ultra-fast matrix multiplication [1, 2, 3, 4, 5]. Typically, the photonic circuits are sensitive to thermal variation as temperature impacts the refractive index of the optical component. Especially for compact microring resonator (MRR)-based photonic accelerators, the weight $w$ is encoded by the differential transmission of the add-drop MRR as $w=g(2a-1);~{}a=\frac{\alpha^{2}-2r\alpha\cos{\phi}+r^{2}}{1-2r\alpha\cos{\phi}+r^{2}\alpha^{2}}\in(0,1)$, where $\alpha$ is the attenuation factor, $r$ is the coupling factor, $\phi$ is the round-trip phase shift, $a$ is the through-port transmission, and $g$ is the scaling factor. To accommodate multiple channels in a single free-spectral range (FSR) while minimizing spectral crosstalk, i.e., large finesse, MRR is designed to have a mid-to-high quality factor. This design choice makes the MRR highly sensitive to small disturbances, highlighting the urgency in addressing challenges related to thermal variation robustness.
@@ II-B Noise-Aware ONN Optimization
Noise-aware optimization for photonic accelerators includes two categories, i.e., offline optimization and on-chip optimization. Prior offline methods inject noise into ONN model training to obtain smooth solution space for better noise tolerance [10]. However, such a method requires accurate noise modeling and can only handle static, known variations, which cannot be adapted to dynamically drifting, unknown variations. On-chip training, as a trend, has been demonstrated for online adaptability and efficiency in in-situ accuracy recovery. The pretrained optical NNs are fine-tuned on a target dataset with on-chip gradient calculation [19, 20, 21]. However, they rely on accurate gradient calculation and require costly forward and backward propagation on a given training set, which can significantly degrade the edge inference throughput and efficiency and bear data privacy issues.
@@ III-B1 Device Phase Variation
Beside nonuniform noise distribution, the noise profiles are dynamically drifting over time, as shown in Fig. 5. We simulate such dynamics with two-level sampling. For each PTC, we use the Standard Deviation ($\sigma_{ij}$, std) of the phase noise distribution to represent the noise intensity for the MRR at coordinate $ij$, while the intensity itself is time-varying. At time step $t$, we sample the step-$t$ noise intensity from a distribution, i.e., $\sigma_{ij}^{t+1^{\prime}}\sim\mathcal{N}(\mu_{s}(t),\sigma_{s}^{2}(t))$. We apply exponential moving average $\sigma_{ij}^{t+1}\leftarrow\beta_{\sigma}\sigma_{ij}^{t}+(1-\beta_{\sigma})\sigma_{ij}^{t+1^{\prime}}$ with a dampling factor $\beta_{\sigma}$ to smooth out the intensity drifting. $\mu_{s}(t)$ and $\sigma_{s}(t)$ are temporally evolving with a certain scheduling function.
@@ III-B2 Environmental Temperature Drift
The environmental temperature fluctuation can impact the reliability of the chip, especially for thermal-sensitive microring resonator-based PTCs [27, 23, 24]. As shown in Fig. 1(b), for MRRs, even 0.05K temperature drift can lead to a relatively large resonant wavelength shift, leading to large errors on the represented weight value. We assume a linear dependence of on-resonance wavelength $\lambda_{c}$ on temperature $T$. Thus, we assume a constant $\partial\lambda/\partial T$. The drift on the round-trip phase shift $\delta\phi$ due to temperature change is derived as,
@@ III-C Salience-Aware Sparse Calibration
(2) Accurate gradients – Unlike on-chip training, where gradient errors will exponentially accumulate through layers, our method only calculates gradients for each block without backpropagation. Hence, our method avoids slow convergence or divergence issues even under large noise.
@@ IV-A Simulation Setup
Dataset and Models.  We evaluate our method on a three-layer CNN (C64K3-C64K3-C64K3-Pool5-FC10) on Fashion-MNIST[31], VGG-8[32] on CIFAR-10[33], and ResNet-18[34] on CIFAR-100[33] for image classification.
@@
Architecture Settings.  As a challenging case study, the hardware platform in this work is assumed to be a multi-tile, multi-core photonic tensor accelerator based on a thermally sensitive MRR weight bank shown in Fig. 3. Note that our method is not specific to MRR weight banks but can generalize to all universal optical matrix-vector multiplication units. We assume that the photonic accelerator has 4 tiles, and each tile has 4 cores. Each core is an 8$\times$8 add-drop MRR weight bank that can perform 8$\times$8 matrix-vector multiplication per core per cycle. Detailed architecture description is in Section III-A. The MRR device modeling is based on Section II-A. The device/circuit variation modeling is based on Section III-B.
@@
(1) Phase Variation (PV). To simulate real-world chip noise scenarios, two cases are created: "Low Noise" and "High Noise," corresponding to low and high standard deviation (std.) in phase shift, modeled as $\Delta\phi\in\mathcal{N}(0,\sigma^{2})$. As explained in Section III-B, the noise std. $\sigma_{ij}^{t+1^{\prime}}$ for each MRR device is time-variant and sampled at each time step from a noise level map $\mathcal{N}\in(\mu_{s}(t),\sigma_{s}^{2}(t))$. The noise level distribution gradually changes its high-noise region from the chip’s left edge to the top-left corner ("Edge-to-Corner") to capture dynamic noise profile drifting. Specific noise level functions are chosen for the Low and High Noise cases:
@@
(2) Temperature Drift (TD): Two cases are designed to represent different types of temperature change with $t_{max}=20,000$:
@@
(3) Thermal Crosstalk (CT):, we use a default MRR spacing $l_{v}=200\mu m$, $l_{h}=60\mu m$, crosstalk coefficient $k_{1}=0.1$, and model crosstalk only among MRRs within the same PTC [28, 25, 30].
@@ IV-C Main Results: Compare with Prior Work
Across all benchmarks, BS1 suffers from severe accuracy degradation (30%-60%). Noise-aware training [10], though it enhances the smoothness of the solution space of the model, shows limited effect on accuracy improvement, as the noise distribution used in pre-training is significantly different from the physical variations. Note that our DOCTOR is orthogonal to noise-aware training, allowing simultaneous application for a solution space that is both locally smooth and adaptive to the drifting noise distribution. On-chip training [21, 22] can boost the accuracy on small networks, e.g., 8% on CNN-FMNIST and VGG8-CIFAR10. However, it performs poorly on deep ResNet18 and is not stable on certain benchmarks, e.g., CT+PV.2+TD.1. The fundamental reason is that the gradient estimation error exponentially accumulates with backpropagation, which can lead to poor training performance and even divergence. Also, though it is only trained for 1 epoch on a 10% training set, it consumes 5-6$\times$ more runtime (cycles) in training, which is not practical to be deployed on throughput-restricted edge platforms. In contrast, our proposed DOCTOR method can stably maintain high accuracy even with rapidly drifting noise distributions and temperature with less than 1-2.5% accuracy drop at the cost of merely 0.1%-5.1% cycle overhead. On average, DOCTOR is +34% more accurate and 2-3 orders-of-magnitude more efficient than on-chip training. We visualize our DOCTOR flow in Fig. 13 to show the temperature drift is detected and the accuracy is rapidly resumed with only 3.6% cycle overhead.